# EMNLP: Educator-role Moral and Normative Large Language Models Profiling

### v2

@@ Appendix H.2 RQ2 LLM Stages of Moral Development at All Temperatures
Table 15: MSS of 14 LLMs Across 5 Moral Dimensions (English, temperature $t\!=\!0\text{\textendash}1$) with Heat-map Coloring and Overall Average (88 dilemmas)
@@
Table 16: MSS of 14 LLMs Across 5 Moral Dimensions (Chinese, temperature $t\!=\!0\text{\textendash}1$) with Heat-map Coloring and Overall Average (88 dilemmas)
@@ Appendix H.3 RQ3 Harmful Response Rates under Soft Prompt Injection
The results in Table 17 present model-level HRR across all flaw types. The analysis reveals marked differences in ethical robustness under role injection, highlighting how certain models—especially those with higher reasoning capacity—are more susceptible to harmful outputs when simulating the teacher role.
@@
Table 17: HRR of LLMs Across 4 Potential Moral Flaw Dimensions (temperature $t\!=\!0\text{\textendash}1$) with Heat-map Coloring and Overall Average (100 items)
Model Temp INC OFF IND IR Overall Reasoning-Oriented Models DeepSeek-R1 0.00 1.00 1.00 0.96 0.71 0.92 0.25 1.00 1.00 1.00 0.40 0.85 0.50 1.00 1.00 1.00 0.40 0.85 0.75 1.00 1.00 1.00 0.44 0.86 1.00 1.00 1.00 1.00 0.48 0.87 Claude-3.7 0.00 1.00 0.60 0.96 0.00 0.64 0.25 1.00 0.56 1.00 0.00 0.64 0.50 1.00 0.60 1.00 0.00 0.65 0.75 1.00 0.48 1.00 0.08 0.64 1.00 1.00 0.56 0.96 0.12 0.66 R1-Distill-70B 0.00 0.64 0.92 1.00 0.20 0.69 0.25 0.72 1.00 0.92 0.24 0.72 0.50 0.52 1.00 1.00 0.16 0.67 0.75 0.44 1.00 0.96 0.36 0.69 1.00 0.48 1.00 1.00 0.16 0.66 QwQ-32B 0.00 0.96 1.00 1.00 0.32 0.82 0.25 1.00 1.00 1.00 0.40 0.85 0.50 0.96 1.00 1.00 0.40 0.84 0.75 1.00 1.00 1.00 0.40 0.85 1.00 1.00 1.00 1.00 0.48 0.87 R1-Distill-8B 0.00 0.48 0.80 0.84 0.08 0.55 0.25 0.32 0.84 0.84 0.16 0.54 0.50 0.52 0.80 0.72 0.12 0.54 0.75 0.24 0.80 0.72 0.08 0.46 1.00 0.16 0.76 0.72 0.08 0.43 GLM-Z1-9B 0.00 0.96 1.00 1.00 0.60 0.89 0.25 0.92 1.00 1.00 0.48 0.85 0.50 0.96 1.00 1.00 0.48 0.86 0.75 0.88 1.00 1.00 0.60 0.87 1.00 1.00 1.00 1.00 0.60 0.90 Qwen3-235B-R 0.00 1.00 1.00 1.00 0.56 0.89 0.25 1.00 1.00 1.00 0.48 0.87 0.50 1.00 1.00 1.00 0.40 0.85 0.75 1.00 1.00 0.96 0.60 0.89 1.00 1.00 1.00 1.00 0.44 0.86 Non-Reasoning Models GPT-4.1 0.00 1.00 1.00 0.84 0.44 0.86 0.25 1.00 1.00 0.88 0.44 0.87 0.50 0.96 1.00 0.92 0.41 0.86 0.75 1.00 1.00 0.88 0.44 0.86 1.00 1.00 1.00 0.92 0.33 0.85 DeepSeek-V3 0.00 1.00 1.00 1.00 0.20 0.80 0.25 1.00 1.00 0.96 0.42 0.85 0.50 1.00 1.00 1.00 0.12 0.78 0.75 1.00 1.00 1.00 0.20 0.80 1.00 1.00 1.00 1.00 0.12 0.78 Qwen2.5-72B 0.00 0.60 0.32 0.52 0.00 0.36 0.25 0.72 0.36 0.52 0.00 0.40 0.50 0.60 0.44 0.60 0.00 0.41 0.75 0.52 0.52 0.60 0.00 0.41 1.00 0.52 0.60 0.48 0.00 0.40 Qwen2.5-32B 0.00 0.36 0.96 0.52 0.00 0.46 0.25 0.32 0.92 0.52 0.00 0.44 0.50 0.20 0.96 0.56 0.00 0.43 0.75 0.32 0.96 0.44 0.00 0.43 1.00 0.36 0.96 0.48 0.00 0.45 Qwen2.5-7B 0.00 0.08 0.52 0.28 0.00 0.22 0.25 0.08 0.48 0.20 0.04 0.20 0.50 0.12 0.48 0.20 0.04 0.21 0.75 0.08 0.40 0.20 0.00 0.17 1.00 0.08 0.44 0.12 0.04 0.17 Baichuan2-7B 0.00 0.28 0.12 0.08 0.04 0.13 0.25 0.32 0.12 0.28 0.12 0.21 0.50 0.12 0.16 0.16 0.04 0.12 0.75 0.16 0.20 0.24 0.00 0.15 1.00 0.16 0.16 0.12 0.00 0.11 Qwen3-235B-C 0.00 0.84 1.00 0.88 0.00 0.68 0.25 0.84 1.00 0.92 0.00 0.69 0.50 0.96 1.00 0.84 0.08 0.72 0.75 0.84 1.00 0.84 0.00 0.67 1.00 0.96 1.00 0.80 0.00 0.69

### v3

@@ Appendix H.2 RQ2 LLM Stages of Moral Development at All Temperatures
Table 15: MSS of 14 LLMs Across 5 Moral Dimensions (English, temperature $t\!=\!0\text{–}1$) with Heat-map Coloring and Overall Average (88 dilemmas)
@@
Table 16: MSS of 14 LLMs Across 5 Moral Dimensions (Chinese, temperature $t\!=\!0\text{–}1$) with Heat-map Coloring and Overall Average (88 dilemmas)
@@ Appendix H.3 RQ3 Harmful Response Rates under Soft Prompt Injection
The results in Table 17 and  18 present model-level HRR across all flaw types. The analysis reveals marked differences in ethical robustness under role injection, highlighting how certain models—especially those with higher reasoning capacity—are more susceptible to harmful outputs when simulating the teacher role.
@@
Table 17: HRR of Reasoning-Oriented LLMs Across 4 Potential Moral Flaw Dimensions (temperature $t\!=\!0\text{–}1$) with Heat-map Coloring and Overall Average (100 items)
Model Temp INC OFF IND IR Overall Reasoning-Oriented Models DeepSeek-R1 0.00 1.00 1.00 0.96 0.71 0.92 0.25 1.00 1.00 1.00 0.40 0.85 0.50 1.00 1.00 1.00 0.40 0.85 0.75 1.00 1.00 1.00 0.44 0.86 1.00 1.00 1.00 1.00 0.48 0.87 Claude-3.7 0.00 1.00 0.60 0.96 0.00 0.64 0.25 1.00 0.56 1.00 0.00 0.64 0.50 1.00 0.60 1.00 0.00 0.65 0.75 1.00 0.48 1.00 0.08 0.64 1.00 1.00 0.56 0.96 0.12 0.66 R1-Distill-70B 0.00 0.64 0.92 1.00 0.20 0.69 0.25 0.72 1.00 0.92 0.24 0.72 0.50 0.52 1.00 1.00 0.16 0.67 0.75 0.44 1.00 0.96 0.36 0.69 1.00 0.48 1.00 1.00 0.16 0.66 QwQ-32B 0.00 0.96 1.00 1.00 0.32 0.82 0.25 1.00 1.00 1.00 0.40 0.85 0.50 0.96 1.00 1.00 0.40 0.84 0.75 1.00 1.00 1.00 0.40 0.85 1.00 1.00 1.00 1.00 0.48 0.87 R1-Distill-8B 0.00 0.48 0.80 0.84 0.08 0.55 0.25 0.32 0.84 0.84 0.16 0.54 0.50 0.52 0.80 0.72 0.12 0.54 0.75 0.24 0.80 0.72 0.08 0.46 1.00 0.16 0.76 0.72 0.08 0.43 GLM-Z1-9B 0.00 0.96 1.00 1.00 0.60 0.89 0.25 0.92 1.00 1.00 0.48 0.85 0.50 0.96 1.00 1.00 0.48 0.86 0.75 0.88 1.00 1.00 0.60 0.87 1.00 1.00 1.00 1.00 0.60 0.90 Qwen3-235B-R 0.00 1.00 1.00 1.00 0.56 0.89 0.25 1.00 1.00 1.00 0.48 0.87 0.50 1.00 1.00 1.00 0.40 0.85 0.75 1.00 1.00 0.96 0.60 0.89 1.00 1.00 1.00 1.00 0.44 0.86
Table 18: HRR of Non-Reasoning LLMs Across 4 Potential Moral Flaw Dimensions (temperature $t\!=\!0\text{–}1$) with Heat-map Coloring and Overall Average (100 items)
Model Temp INC OFF IND IR Overall Non-Reasoning Models GPT-4.1 0.00 1.00 1.00 0.84 0.44 0.86 0.25 1.00 1.00 0.88 0.44 0.87 0.50 0.96 1.00 0.92 0.41 0.86 0.75 1.00 1.00 0.88 0.44 0.86 1.00 1.00 1.00 0.92 0.33 0.85 DeepSeek-V3 0.00 1.00 1.00 1.00 0.20 0.80 0.25 1.00 1.00 0.96 0.42 0.85 0.50 1.00 1.00 1.00 0.12 0.78 0.75 1.00 1.00 1.00 0.20 0.80 1.00 1.00 1.00 1.00 0.12 0.78 Qwen2.5-72B 0.00 0.60 0.32 0.52 0.00 0.36 0.25 0.72 0.36 0.52 0.00 0.40 0.50 0.60 0.44 0.60 0.00 0.41 0.75 0.52 0.52 0.60 0.00 0.41 1.00 0.52 0.60 0.48 0.00 0.40 Qwen2.5-32B 0.00 0.36 0.96 0.52 0.00 0.46 0.25 0.32 0.92 0.52 0.00 0.44 0.50 0.20 0.96 0.56 0.00 0.43 0.75 0.32 0.96 0.44 0.00 0.43 1.00 0.36 0.96 0.48 0.00 0.45 Qwen2.5-7B 0.00 0.08 0.52 0.28 0.00 0.22 0.25 0.08 0.48 0.20 0.04 0.20 0.50 0.12 0.48 0.20 0.04 0.21 0.75 0.08 0.40 0.20 0.00 0.17 1.00 0.08 0.44 0.12 0.04 0.17 Baichuan2-7B 0.00 0.28 0.12 0.08 0.04 0.13 0.25 0.32 0.12 0.28 0.12 0.21 0.50 0.12 0.16 0.16 0.04 0.12 0.75 0.16 0.20 0.24 0.00 0.15 1.00 0.16 0.16 0.12 0.00 0.11 Qwen3-235B-C 0.00 0.84 1.00 0.88 0.00 0.68 0.25 0.84 1.00 0.92 0.00 0.69 0.50 0.96 1.00 0.84 0.08 0.72 0.75 0.84 1.00 0.84 0.00 0.67 1.00 0.96 1.00 0.80 0.00 0.69